OpenStudy (anonymous):

To the nearest tenth, what is the distance between points (7,4) and (12,5)? Please explain how you got the answer!

OpenStudy (anonymous):

\[\sqrt{(12-7)^{2}+(5-4)^{2}}=\sqrt{25+1}=\sqrt{26}\]